MOLECULAR SPECTROSCOPIC STUDIES OF METAL(II)[Mn(II),Co(II) AND Ni(II)] HALOGEN COM-PLEXES WITH METHYLANILINES

\underline{M. KUMRU}, Fatih University, Faculty of Arts and Sciences, Department of Physics, 34500 Büyükcekmece, Istanbul, Turkey; K. GOLCUK, Department of Chemistry, University of Michigan, 930 N. University, Ann Arbor, MI 48109-1055; A. ALTUN, Max-Planck-Institut für Kohlenforschung, Kaiser-Wilhelm-Platz 1, 45470 Mülheim an der Ruhr, Germany.